OpenStudy (anonymous):

Write the equation in standard form of the line that passes through thepoints and has the given slope Points= (2,-3) slope= -4

OpenStudy (anonymous):

\[y=-4x+5\]

OpenStudy (anonymous):

IT has to be standard form

OpenStudy (anonymous):

\[y-y _{1}=m(x-x_1)\] Substitute coordinates: \[y-(-3)=(-4)(x-2)\] And then simplify equation.

OpenStudy (anonymous):

\[ y+3 = -4x +8 \Rightarrow y=−4x+5 \Rightarrow 4x+y=5 \]